how do i solve 9(20-12t)=8(6+9t)-24
i keep getting different answers.

Answer by edjones(8007)

9(20-12t)=8(6+9t)-24
180-108t=48+72t-24
180-108t=72t+24
-180t=-156
t=13/15
.
Ed